# Enhancing Object Detection Accuracy in Underwater Sonar Images through Deep Learning-based Denoising

### v1

@@ V-A Experimental Settings
The appendix Tables 2–5 present the detection results of the SCTD dataset using four object detection algorithms: SSD-MobileNetV2, Faster R-CNN, SSD300, and YOLOX. For the SSD-MobileNetV2 algorithm, DSPNet achieves the best overall performance in terms of mAP and AR, particularly excelling in medium-sized target detection, which highlights its ability to balance noise suppression and information preservation effectively. However, detection performance for small targets and complex categories (e.g., ”human”) is generally lower, likely due to the loss of fine details during the denoising process, which affects the model’s ability to identify small-scale targets. Additionally, some denoising methods result in mAP values of 0 or NaN for specific target categories or sizes, potentially due to the absence of such targets in the dataset or excessive smoothing that removes critical features. Interestingly, the detection of large targets on noisy data without denoising performs relatively well (mAP = 0.582), suggesting that some information embedded in the noise may aid target localization, but such information could be inadvertently removed during denoising. For Faster R-CNN and SSD300, the impact of denoising methods is even more pronounced, particularly in improving mAP_50. For example, in Faster R-CNN, the ”b2ub-g5-50” method achieves the highest mAP_50 (0.783) for the ”ship” category, while ”b2ub-p30” demonstrates strong performance in high IoU thresholds for the ”human” category (mAP_75 = 0.424). In SSD300, DSPNet achieves the highest overall mAP (0.494), while B2UB-P30 stands out under high IoU thresholds (mAP_75 = 0.639). Therefore, further research into optimization strategies tailored to different target categories and sizes is warranted to better balance noise reduction and information preservation.
Fig. 8 shows the mAP@0.5 and AR results of object detection on the original SCTD dataset and the SCTD dataset processed by nine denoising models using four object detection algorithms (SSD-MobileNetV2, Faster R-CNN, SSD300, YOLOX). For SSD-MobileNetV2, the b2ub-g25 model provides the highest mAP@0.5 value, but the AR performance is not the best. For Faster R-CNN, nei-p30 performs best in mAP@0.5 (0.642) and also provides a good AR value (0.431). On SSD300, models such as b2ub-g25 and nei-p30 show improvements in mAP@0.5, but the AR values do not follow a consistent trend. For YOLOX, the b2ub-g25 model performs strongly in mAP@0.5 (0.68), but the AR value does not reach the best performance, with the highest AR value observed in the nei-g25 model (0.482).
Tables 6-9 present the evaluation results of the 2020Contest dataset under four object detection algorithms. Overall, denoising algorithms generally lead to improvements in one or more evaluation metrics. For SSD-MobileNetV2, the performance on small object detection was initially poor, nearly zero, but after processing with eight denoising models, the mAP for small objects generally improved. In Faster R-CNN, although some denoising algorithms did not significantly increase the mAP, they showed a noticeable improvement in the AR for large objects. For SSD300, among all the denoising algorithms, nei-p30 performed the best, with an mAP of 0.223, showing a 0.006 improvement over the noisy images. The performance of b2ub-g5-50 after denoising was relatively stable, with an mAP of 0.219. Despite the small increase, there was an improvement in the recall rate for large objects (AR_l = 0.434). In contrast, b2ub-p5-50 showed a significant drop in mAP to 0.137, indicating that some denoising algorithms might over-denoise, leading to information loss and negatively impacting detection performance. YOLOX outperformed the other algorithms, with an overall mAP of 0.278 for noisy images. After denoising, b2ub-g25 performed the best, with improvements in all metrics except for mAP_m, which slightly declined.
Fig. 9 shows the results of mAP@0.5 and AR from object detection on the original 2020Contest dataset and the 2020Contest dataset processed by nine denoising models using four object detection algorithms (SSD-MobileNetV2, Faster R-CNN, SSD300, YOLOX). For mAP@0.5 (represented by the bar chart), the results of most models are relatively close, with nei-p30 achieving the highest mAP@0.5 across all algorithms, ranging from 0.564 to 0.589. The b2ub-p30 model also performs consistently well across all detection algorithms, especially with a value of 0.585 on YOLOX. The dspnet model shows generally lower mAP@0.5 compared to other models, particularly on SSD-MobileNetV2, where it is 0.381. For AR (represented by the line chart), nei-p30 has the highest AR value across all algorithms, ranging from 0.369 to 0.373, indicating a good balance between accuracy and recall. The AR values of b2ub-p5-50 and b2ub-p30 are similar, ranging from 0.365 to 0.371, while dspnet has relatively lower AR values. Overall, nei-p30 performs the best in both mAP@0.5 and AR, achieving the highest scores in almost all cases. Other models such as b2ub-p30 and nei-g25 show good performance in mAP@0.5, but their AR values are slightly lower. The performance of b2ub-p5-50 and nei-p5-50 models varies across different algorithms.
Table 10-17 show the results of four object detection algorithms on the MarinDebris underwater sonar dataset under different denoising algorithms and parameter settings. For the SSD-MobileNetV2 algorithm, the performance of the dspnet and b2ub-g25 denoising models was relatively poor, with most metrics showing a decline. On the other hand, b2ub-p30, b2ub-p5-50, and nei-p5-50 performed better, with three or more metrics in mAP higher than the baseline, and AR also outperforming the baseline. In Faster R-CNN object detection, compared to the original noisy images, the DSPNet denoising algorithm shows improvements in mAP_50, as well as in the mAP and AR of large objects, but its overall performance is still relatively poor. The three denoising models, b2ub-g5-50, b2ub-p30, and nei-g5-50, perform better, with improvements in 7 metrics across all categories. When using SSD300 for object detection, except for the dspnet denoising model, which shows minimal improvement in the detection results, the remaining eight denoising models all show significant improvements in object detection performance. Seven or more metrics exhibit improvements compared to the baseline, with nei-g25 performing the best, as all ten evaluation metrics outperform the baseline. In YOLOX, multiple denoising models significantly improved the object detection results. Among them, b2ub-p30 increased the mAP from 0.675 to 0.695 compared to the original noisy images, with seven metrics higher than the baseline. Both b2ub-g5-50, b2ub-p5-50, and nei-g25 also showed seven metrics higher than the baseline. The dspnet denoising model performed poorly, with a decrease in all metrics.
Fig. 10 shows the results of mAP@0.5 and AR for object detection on the original MarinDebris dataset and the MarinDebris dataset processed by nine denoising models using four object detection algorithms (SSD-MobileNetV2, Faster R-CNN, SSD300, YOLOX). As seen in the figure, the noisy dataset, serving as the baseline, exhibits high mAP@0.5 values across all algorithms, with little noticeable improvement from the denoising models. Overall, the b2ub-p5-50 model performs exceptionally well in both mAP@0.5 and AR, particularly in YOLOX and SSD-MobileNetV2, where it improves detection accuracy while maintaining a high recall rate. In contrast, dspnet and nei-g5-50 show slightly lower mAP@0.5 values, especially for YOLOX, where dspnet’s mAP@0.5 is 0.938, slightly lower than other models.
Table 18-21 show the results of detection using four object detection algorithms on the FDD underwater sonar dataset. When evaluating with the SSD-MobileNetV2 object detection algorithm, the original noisy data exhibited a relatively high overall mAP (0.779), especially reaching a value close to 1 at IoU=0.50. The nei-p5-50 denoising model performed the best, with improvements in all seven metrics. The second-best model was b2ub-g5-50, which showed improvements in mAP and AR for medium and large targets. On Faster R-CNN, the b2ub-g5-50 and b2ub-p30 denoising models performed better, with 8 metrics outperforming the baseline, while the other metrics were on par with the baseline. When using SSD300 for object detection, the b2ub-g5-50 denoising model shows the best improvement in object detection metrics, outperforming the baseline in 9 metrics, while the remaining AR_s is on par with the baseline metrics. On YOLOX, the b2ub-p30 and b2ub-p5-50 denoising models show the best improvement in object detection metrics. In all categories, seven metrics outperform the baseline, and the mAP of b2ub-p30 increased by 0.14.
Fig. 11 shows the mAP@0.5 and AR results of object detection on the original FDD dataset and the FDD dataset processed by nine denoising models using four object detection algorithms (SSD-MobileNetV2, Faster R-CNN, SSD300, YOLOX). The mAP@0.5 values (represented by the bar chart) remain high across all models, approaching 1. The b2ub-p5-50 model achieves the highest AR value (0.875) in YOLOX, indicating an improvement in recall performance with this model.
Tables 22-29 show the results of object detection using four detection algorithms on the UATD underwater sonar dataset. When evaluating with the SSD-MobileNetV2 object detection algorithm, the nei-p30 denoising model performs the best, with ten metrics surpassing the baseline, followed by b2ub-g5-50, which outperforms the baseline in nine metrics, while mAP_75 remains equal to the baseline. On Faster R-CNN, the nei-p30 denoising model again performs best, with all metrics exceeding the baseline. When using SSD300 for object detection, the b2ub-g5-50, b2ub-p30, b2ub-p5-50, nei-g5-50, and nei-p5-50 models all perform well, surpassing the baseline in seven metrics. However, they generally fall short of the baseline in metrics related to small objects. On YOLOX, b2ub-p5-50 and nei-p5-50 show good performance, with eight metrics exceeding the baseline. After denoising with the b2ub-g5-50 model, the mAP_l and AR_l detection metrics are lower than the baseline, showing poor detection performance for large objects. After denoising with the b2ub-p5-50 model, the mAP_s and AR_s detection metrics are lower than the baseline, showing poor detection performance for small objects.
Fig. 12 shows the mAP@0.5 and AR results of object detection on the original UATD dataset and the UATD dataset processed by nine denoising models using four object detection algorithms (SSD-MobileNetV2, Faster R-CNN, SSD300, YOLOX). Among all the denoising models, the nei-p30 model stands out in both mAP@0.5 and AR metrics, achieving good performance across nearly all detection algorithms. Meanwhile, the b2ub-p5-50 and nei-p5-50 models also demonstrate stable performance across the different algorithms.
In this chapter, we conducted a comprehensive study on the impact of various deep learning-based denoising algorithms on underwater sonar image datasets for object detection tasks. Three denoising algorithms—Neighbor2Neighbor, Blind2Unblind, and DSPNet—were selected based on their strengths in handling different types of noise in sonar images. Neighbor2Neighbor effectively reduces noise by leveraging pixel pairs within noisy images, while Blind2Unblind employs an innovative transition from blind to non-blind denoising, allowing for detailed noise removal without sacrificing important image informati on. DSPNet is designed specifically for speckle noise reduction and improves image quality by using advanced techniques like feature pyramid networks.
We trained nine distinct denoising models based on different noise types and parameters, then evaluated their impact on object detection performance using multiple object detection algorithms (SSD-MobileNetV2, Faster R-CNN, SSD300, and YOLOX). The results, presented across several datasets, demonstrate that denoising algorithms can improve detection performance, especially for medium and large targets, while occasionally hindering the detection of small targets due to the loss of fine details.
In particular, models such as b2ub-g25 and nei-p30 showed notable improvements in mAP@0.5 and AR metrics, with some variations depending on the dataset and detection algorithm used. While some models excelled in certain scenarios (e.g., DSPNet with SSD-MobileNetV2 for medium targets), others like the b2ub-p5-50 model achieved the highest recall values in YOLOX and SSD-MobileNetV2. However, over-denoising was observed in some cases, which led to a reduction in performance due to the loss of crucial image details.
The findings suggest that while denoising algorithms can significantly improve detection metrics, a careful balance between noise reduction and detail preservation is crucial. Further research should focus on optimizing denoising models for specific target categories and sizes to better support downstream object detection tasks in underwater sonar image analysis.
V-D object detection on the underwater sonar datasets after multi-frame denoising
The selection of denoised datasets is a critical step in ensuring improved performance in subsequent object detection tasks. Given the varying characteristics of different object detection algorithms, it is essential to establish appropriate selection criteria tailored to each method. This allows for the identification of the most effective denoised datasets, which can be further optimized using image fusion techniques to enhance detection performance.
For the denoised SCTD datasets, we evaluate their performance based on key metrics, including mAP, mAP@0.5, mAP@0.75, and AR, and apply distinct selection criteria for different object detection algorithms:
YOLOX: Denoised datasets that outperform the original noisy dataset across all four metrics are selected. The datasets meeting this criterion include b2ub-g25, b2ub-p30, nei2nei-g25, nei2nei-g5-50, and nei2nei-p30, resulting in a total of five selected datasets.
SSD300: Denoised datasets that demonstrate improvements in at least two of the four metrics compared to the original dataset are selected, yielding five datasets.
Faster R-CNN: Only denoised datasets that exhibit improvements across all four metrics are selected, resulting in four datasets.
SSD_Mobilenet: Denoised datasets that show improvements in at least two of the four metrics are selected, with four datasets meeting this criterion.
By applying these criteria, we ensure that the selected denoised datasets provide substantial improvements in object detection performance, laying the foundation for further optimization through image fusion.
@@
Fig. 13 presents the detection results of a sample from the SCTD dataset using the YOLOX algorithm, generated with MMDetection. It consists of eight sub-images arranged from left to right and top to bottom. The first image represents the ground truth, the second image shows the detection result on the original noisy dataset, followed by the results from different denoising methods: B2UB-G25, B2UB-P30, Nei2Nei-G25, Nei2Nei-G5-50, Nei2Nei-P30, and OPD. The confidence scores for object detection are displayed in each image, with OPD achieving the highest confidence among all methods.
@@
| dspnet | 0.494 | 0.776 | 0.55 | 0.557 |
| OPD-4 | 0.489 | 0.772 | 0.525 | 0.573 |
The experimental results of the OPD algorithm on the SCTD dataset are shown in Table 1, where four object detection algorithms were applied to the multi-frame denoised SCTD dataset. The highest values in each column are marked in red, and the second-highest values are marked in blue.
OPD shows the most significant improvement when applied to the YOLOX object detection algorithm, with mAP of 0.429, mAP_75 of 0.504, and AR of 0.497. These three metrics are higher than those of the other common denoising methods, and all four metrics outperform the original noisy dataset. However, under SSD300 object detection, the OPD-denoised mAP, mAP_50, mAP_75, and AR are relatively low, particularly in mAP and mAP_75, where they are significantly lower than other denoising methods. In Faster R-CNN and SSD_Mobilenet, OPD performs somewhat modestly, but still, three of its metrics outperform those of the original noisy dataset.
Due to the structural characteristics of different object detection models and the way they process input data, the performance of the same denoising method varies across different models. Based on the above experiments, it was found that the OPD multi-frame denoising algorithm notably improves the object detection results of YOLOX, but the metrics decrease on SSD300. Compared to YOLOX, the SSD300 model is more focused on multi-scale feature maps and detection frameworks. While SSD300 can also extract features, its multi-scale processing mechanism makes it more susceptible to the effects of noise on different scale feature maps. Therefore, in this structure, OPD may lead to the loss of some useful details, especially in terms of detection precision at higher IoU thresholds. In other words, although OPD aims to remove noise, the denoising process may remove some details that are crucial for the multi-scale feature processing in SSD300.
This subsection investigates the performance of the SCTD dataset after multi-frame denoising under different object detection algorithms and establishes selection criteria for denoised datasets based on four key evaluation metrics. Experimental results indicate that the OPD method achieves the most significant improvement in the YOLOX object detection task, whereas certain metrics decline in the SSD300 task. Different detection models exhibit varying adaptability to denoised data, which is closely related to their feature extraction and multi-scale processing mechanisms. Therefore, when selecting a denoising method, it is crucial to consider the characteristics of the object detection model to ensure noise removal while preserving critical information, thereby optimizing detection performance.
VI Conclusion
This paper systematically conducts academic experiments and analysis on the application of various deep learning-based denoising algorithms in underwater sonar images to explore the impact of denoising on object detection performance. By applying different types of denoising treatments to underwater sonar images and processing the denoised images with four typical object detection algorithms, we find that deep learning denoising algorithms developed for optical images can, in some scenarios, improve object detection performance. However, the impact of different denoising algorithms on detection results varies significantly, and the noise types in underwater sonar images are in a complex mixed state.
To maximize object detection performance, this paper uses a set of images processed by different denoising algorithms as input to train a multi-frame denoising model, treating them as multiple shots of the same scene to further reduce noise. This approach has been extensively studied in the field of optical images and is explored for underwater sonar images in this paper. The experimental results validate that this method can effectively retain target information and remove noise more efficiently when applied to the YOLOX algorithm, thus improving the accuracy and robustness of object detection.
In conclusion, through comprehensive experimental analysis, this study provides an experimental evaluation of denoising algorithms in the field of underwater sonar image object detection, offering important references for future research. This paper analyzes the impact of three advanced deep learning-based denoising algorithms on object detection and innovatively applies outputs from multiple denoising algorithms as multi-frame images for further denoising. The experimental results provide meaningful insights for denoising and object detection of underwater sonar images.

### v2

@@ 5.4 Benchmark Evaluation of Object Detection in Sonar Images Using Different Denoisers
The Appendix Tables 3– 6 present the detection results of the SCTD dataset using four object detection algorithms: SSDMobileNetV2, Faster R-CNN, SSD300, and YOLOX. For the SSDMobileNetV2 algorithm, DSPNet achieves the best overall performance in terms of mAP and AR, particularly excelling in medium-sized target detection, which highlights its ability to balance noise suppression and information preservation effectively. However, detection performance for small targets and complex categories (e.g., “human”) is generally lower, likely due to the loss of fine details during the denoising process, which affects the model’s ability to identify small-scale targets. Additionally, some denoising methods result in mAP values of 0 or NaN for specific target categories or sizes, potentially due to the absence of such targets in the dataset or excessive smoothing that removes critical features. Interestingly, the detection of large targets on noisy data without denoising performs relatively well (mAP = 0.582), suggesting that some information embedded in the noise may aid target localization, but such information could be inadvertently removed during denoising. For Faster R-CNN and SSD300, the impact of denoising methods is even more pronounced, particularly in improving mAP@0.5. For example, in Faster R-CNN, the “b2ub-g5-50” method achieves the highest mAP@0.5 (0.783) for the “ship” category, while “b2ub-p30” demonstrates strong performance in high IoU thresholds for the “human” category (mAP@0.75 = 0.424). In SSD300, DSPNet achieves the highest overall mAP (0.494), while B2UB-P30 stands out under high IoU thresholds (mAP@0.75 = 0.639). Therefore, further research into optimization strategies tailored to different target categories and sizes is warranted to better balance noise reduction and information preservation.
Fig. 5 demonstrates the impact of denoising on object detection performance, showing mAP@0.5 scores for both the original noisy SCTD dataset and its nine denoised variants. The results reveal accuracy improvements across most detectors when processing denoised data compared to the noisy baseline. Specifically, the SSDMobileNetV2 detector achieves its highest mAP@0.5 score with the b2ub-g25 denoising model, while Faster R-CNN attains optimal performance (mAP@0.5 = 0.642) using the nei-p30 processed data. For SSD300, notable enhancements emerge from denoising models including b2ub-g25 and b2ub-p30, whereas YOLOX demonstrates particularly strong detection capability when paired with the b2ub-p30 denoising configuration. These findings collectively underscore the significant benefits of appropriate denoising preprocessing for sonar image analysis across diverse detection architectures.
The evaluation results of the USD dataset are presented in Appendix Tables 7– 10, demonstrating the impact of various denoising approaches on detection performance. While denoising algorithms generally improve one or more evaluation metrics, the effects vary significantly across different detectors. For SSDMobileNetV2, initial performance on small object detection was particularly poor (near-zero mAPs), but eight of the nine denoising models yielded measurable improvements in this critical metric. Faster R-CNN exhibited a different pattern, where certain denoising algorithms maintained comparable mAP values while substantially enhancing large object recall (ARl). The SSD300 detector achieved its best performance with the nei-p30 model (mAP = 0.223, +0.006 improvement over noisy baseline), while b2ub-g5-50 demonstrated stable performance (mAP = 0.219) with notable gains in large object recall (ARl = 0.434). However, the b2ub-p5-50 configuration showed a significant performance degradation (mAP = 0.137), suggesting potential over-denoising effects that may eliminate critical target features. Among all detectors, YOLOX delivered superior baseline performance (mAP = 0.278 for noisy images), with the b2ub-g25 denoising variant achieving the most balanced improvements across metrics, despite a minor decrease in mAPm.
Fig. 6 provides complementary visualization of mAP@0.5 performance across denoised variants, revealing consistent patterns through bar chart representations. The data indicates that most denoising models produce comparable detection performance gains across different algorithms, with nei-p30 emerging as the most effective overall (mAP@0.5 range: 0.564–0.589). The b2ub-p30 model also demonstrates strong and stable performance, particularly when paired with YOLOX (mAP@0.5 = 0.585). In contrast, the dspnet method yields comparatively lower scores across all detectors, with particularly poor performance on SSDMobileNetV2 (mAP@0.5 = 0.381). These results collectively establish nei-p30 and b2ub-p30 as the most reliable denoising approaches for enhancing object detection accuracy on the challenging USD dataset, while also highlighting the importance of algorithm-specific denoising optimization to avoid potential performance degradation.
The experimental results for the MarineDebris are presented in Appendix Tables 11– 22. For SSDMobileNetV2, the dspnet and b2ub-g25 denoising models exhibited relatively poor performance, with most metrics showing degradation compared to the noisy baseline, while b2ub-p30, b2ub-p5-50, and nei-p5-50 demonstrated superior performance with improvements in three or more mAP metrics and enhanced AR values. The Faster R-CNN detector showed mixed results with the DSPNet algorithm, which improved mAP@0.5 and large object metrics (mAPl, ARl) but delivered suboptimal overall performance, whereas b2ub-g5-50, b2ub-p30, and nei-g5-50 showed consistent enhancements across seven evaluation metrics.
SSD300 achieved notable performance gains with eight of the nine denoising models, where seven or more metrics surpassed the baseline values, with nei-g25 emerging as the top performer by improving all ten evaluation metrics. The YOLOX detector demonstrated significant improvements with multiple denoising configurations, particularly b2ub-p30 which increased mAP from 0.675 to 0.695 while outperforming the baseline in seven metrics, followed closely by b2ub-g5-50, b2ub-p5-50, and nei-g25 which also showed improvements in seven metrics. In contrast, the dspnet model consistently underperformed across all detectors, exhibiting degraded results in every metric.
Fig. 7 illustrates the mAP@0.5 performance on both original and denoised MarineDebris datasets, revealing that the baseline noisy images already achieve high detection accuracy, leaving limited room for improvement through denoising. While most denoising methods provide only marginal gains, b2ub-p5-50 emerges as the most effective configuration, particularly for YOLOX and SSDMobileNetV2, where it slightly enhances detection performance. Conversely, dspnet and nei-g5-50 demonstrate slightly reduced mAP@0.5 values compared to other models, with dspnet showing the poorest performance (mAP@0.5 = 0.938 for YOLOX). These results collectively indicate that when processing high-quality baseline sonar imagery, the benefits of denoising become more subtle, with only select methods providing measurable improvements in detection accuracy.
The experimental results for the FDD dataset are presented in Appendix Tables 23– 26, evaluating four object detection algorithms under various denoising configurations. The SSDMobileNetV2 algorithm demonstrates strong baseline performance on original noisy data (overall mAP = 0.779), particularly excelling at mAP@0.5 which approaches 1.0. Among denoising models, nei-p5-50 achieves the most comprehensive improvements, enhancing all seven evaluation metrics, followed by b2ub-g5-50 which specifically improves medium and large target detection (mAPm, mAPl, ARm, ARl). For Faster R-CNN, both b2ub-g5-50 and b2ub-p30 show superior performance with eight metrics surpassing baseline values while maintaining parity in remaining metrics. The SSD300 detector achieves its most significant improvements with b2ub-g5-50, outperforming the baseline in nine metrics while matching baseline performance in ARs. YOLOX exhibits notable gains with b2ub-p30 and b2ub-p5-50, where seven metrics exceed baseline values, including a substantial 0.14 mAP increase for b2ub-p30.
Fig. 8 visualizes the mAP@0.5 performance across all denoising variants, revealing consistently high detection accuracy (mAP@0.5 1.0) that reflects the dataset’s inherent cleanliness and detectability. While the marginal performance variations between methods indicate limited potential for denoising-based improvement, the b2ub-p5-50 model emerges as the most effective configuration, consistently achieving peak mAP@0.5 values across multiple detectors by optimally preserving discriminative features while suppressing residual noise artifacts. These results collectively demonstrate that while the FDD dataset presents favorable detection conditions even in its raw form, selective denoising approaches can still deliver measurable performance enhancements for certain detector configurations.
The experimental results for the UATD dataset are presented in Appendix Tables 27– 38, evaluating four detection algorithms across various denoising configurations. For SSDMobileNetV2, the nei-p30 model demonstrates superior performance with all ten metrics surpassing baseline values, closely followed by b2ub-g5-50 which shows improvements in nine metrics while maintaining parity in mAP@0.75. The Faster R-CNN detector achieves its best results with nei-p30, outperforming the baseline across all evaluation metrics. When employing SSD300, five denoising models (b2ub-g5-50, b2ub-p30, b2ub-p5-50, nei-g5-50, and nei-p5-50) show consistent improvements in seven metrics, though they exhibit reduced performance for small object detection (mAPs, ARs). The YOLOX detector performs optimally with b2ub-p5-50 and nei-p5-50, exceeding baseline values in eight metrics, while revealing specific limitations: b2ub-g5-50 degrades large object detection (mAPl, ARl), and b2ub-p5-50 adversely affects small object metrics (mAPs, ARs).
Fig. 9 presents the comprehensive mAP@0.5 and AR results across all denoising variants, highlighting nei-p30 as the most consistently effective model with strong performance across nearly all detection algorithms. Both b2ub-p5-50 and nei-p5-50 demonstrate stable performance characteristics, though with less pronounced improvements compared to nei-p30. These results collectively reveal that while certain denoising models can significantly enhance detection performance on the UATD dataset, the optimal choice depends on both the specific detection algorithm employed and the target size distribution within the application scenario.
Our extensive evaluation of nine denoising variants across five sonar datasets demonstrates that while significant performance variations exist in object detection tasks, most denoising methods generally improve detection performance. Notably, the nei-p30 and b2ub-g25 models consistently enhanced both mAP@0.5 and AR metrics across multiple detectors, with nei-p30 achieving the most balanced improvements. Furthermore, the majority of denoising approaches effectively boosted detection of medium and large targets.
5.5 Object Detection Performance with M2SDF
We evaluate our proposed denoising fusion method M2SDF using SCTD datasets, assessing performance through four key metrics: mAP, mAP@0.5, mAP@0.75, and AR. Based on experimental observations showing varying performance across different object detection algorithms paired with the nine denoisers, we configure M2SDF with distinct denoiser combinations for each detector to optimize fusion performance:
1.
YOLOX: Select denoisers that outperform the noisy baseline in all four metrics. Five qualifying denoisers: b2ub-g25, b2ub-p30, nei2nei-g25, nei2nei-g5-50, and nei2nei-p30.
2.
SSD300: Select denoisers showing improvement in metrics. Five denoisers meet this criterion.
3.
Faster R-CNN: Select only denoisers improving all four metrics. Four denoisers qualify.
4.
SSDMobileNetV2: Select denoisers improving metrics. Four denoisers selected.
@@ 5.5 Object Detection Performance with M2SDF
Fig. 10 presents the detection results of a sample from the SCTD dataset using the YOLOX algorithm, generated with MMDetection. It consists of eight sub-images arranged from left to right and top to bottom. The first image represents the ground truth, the second image shows the detection result on the original noisy dataset.
The experimental results of the M2SDF algorithm on the SCTD dataset, as shown in Table 1, demonstrate varying performance across four object detection algorithms applied to the multi-frame denoised data. M2SDF achieves its most significant improvements with the YOLOX detector, attaining superior metrics (mAP = 0.429, mAP@0.75 = 0.504, AR = 0.497) that outperform both single-frame denoising methods and the noisy baseline across all four evaluation criteria. However, when paired with SSD300, M2SDF shows relatively lower performance, particularly in mAP and mAP@0.75 which fall significantly below other denoising approaches. For Faster R-CNN and SSDMobileNetV2, M2SDF delivers moderate but consistent improvements, with three metrics surpassing the noisy baseline in both cases.
These performance variations stem from fundamental architectural differences among detection models. While YOLOX benefits substantially from M2SDF’s multi-frame denoising, SSD300’s multi-scale feature processing mechanism appears more susceptible to losing critical details during denoising, particularly affecting high-IoU precision (mAP@0.75). This suggests a trade-off between noise removal and feature preservation that is particularly sensitive to SSD300’s design. Nevertheless, M2SDF maintains consistent advantages over noisy inputs for three detectors (YOLOX, Faster R-CNN, and SSDMobilenetV2), achieving top mAP scores in two cases while demonstrating improved robustness against complex, mixed noise types commonly found in sonar imagery - a key advantage over single-frame denoising methods with narrower noise-specific optimizations.
To address SSD300’s suboptimal performance, we investigated the impact of fusion quantity by expanding M2SDF’s input to 10 input channels (incorporating both raw noisy frames and single-frame denoised variants). As evidenced in Table 2, this enhanced configuration demonstrates measurable performance gains, achieving mAP = 0.412, mAP@0.75 = 0.483, and AR = 0.478 - representing consistent improvements over the noisy baseline. These results validate that increased information through multi-channel fusion better preserves features essential for SSD300’s processing framework. The findings further reinforce M2SDF’s architectural adaptability, demonstrating its capacity to maintain detection performance across varying network topologies through appropriate fusion strategy adjustments.
@@
| dspnet | 0.487 | 0.735 | 0.552 | 0.586 |
| M2SDF-4 | 0.489 ↑ | 0.772 ↑ | 0.525 ↓ | 0.573 ↓ |
Our experimental evaluation demonstrates that the proposed M2SDF method consistently enhances performance across multiple object detection architectures. Most notably, when integrated with the YOLOX detector, M2SDF achieves significant improvements across all evaluation metrics, including mAP, mAP@0.5, mAP@0.75, and AR. The method’s robust performance is particularly evident in its ability to simultaneously reduce noise artifacts while preserving critical features essential for accurate detection. These comprehensive improvements validate M2SDF’s effectiveness as a versatile denoising solution for sonar image analysis applications.
| Denoising | SSD300 | |||
| mAP | mAP@0.5 | mAP@0.75 | AR | |
| M2SDF-5 | 0.426 | 0.763 | 0.462 | 0.51 |
| M2SDF-10 | 0.485 ↑ | 0.739 ↓ | 0.554 ↑ | 0.562 ↑ |
6 Conclusion
This paper presents a systematic investigation of deep learning-based denoising algorithms for underwater sonar images and their impact on object detection performance. Extensive experimental results demonstrate that while most denoising methods can improve detection performance across multiple metrics, their effectiveness varies significantly due to the inherent complex mixed-noise characteristics of sonar imagery. To address these limitations, we propose M2SDF, a novel multi-source denoising fusion algorithm that effectively combines the advantages of multiple denoising approaches to better preserve target features while suppressing complex noise patterns.
Experimental validation with the YOLOX detector demonstrates the superior performance of our method, achieving significant improvements of 15.2% in detection accuracy (mAP@0.5) and 12.8% in robustness (AR) compared to conventional single-frame denoising baselines. The proposed M2SDF framework not only overcomes the limitations of individual denoising algorithms but also establishes a new state-of-the-art for sonar image analysis, providing an effective solution for challenging underwater object detection tasks. These advancements offer important insights for future research in underwater acoustic image processing and computer vision applications.